OpenStudy (anonymous):

Can someone teach me how to do this please?? Medal and fan to best help!! What is the product of sqrt(t)(5 sqrt(3) - 4)? Simplify if possible.

OpenStudy (anonymous):

@dmezzullo ??

OpenStudy (anonymous):

\[\sqrt{5}\sqrt{3}-4, simplifies, \to [\sqrt{15}-4] since \sqrt{5}\sqrt{3}, are, under, common, roots\]

OpenStudy (anonymous):

sorry for the awkward format but that's the only way I could put in all the symbols >.<